OpenStudy (anonymous):

The solution set for the inequality x + 4 (x - 3) includes -11 as an element. True False

OpenStudy (anonymous):

x+4(x-3)=x+4x-12=5x-12 ??

OpenStudy (anonymous):

is it true or false

OpenStudy (anonymous):

i think it's false

OpenStudy (anonymous):

we can write it in this form: 5x-10-2=5(x-2)-2 5(x-2)-2=0 5(x-2)=2 x-2=2/5 x=2/5+2 so we have : x=12/5